OpenStudy (anonymous):

A video game sets the points needed to reach the next level based on the function g(x) = 7(3)x − 1, where x is the current level. The hardest setting promises to multiply the points needed in each level according to the function h(x) = 4x. How many points will a player need on the hardest setting of level 4? a) 16 points b) 189 points c) 3024 points d) 12096 points

OpenStudy (anonymous):

is 3(x)=3^x

OpenStudy (anonymous):

if its score is given by h(x)*g(x) put x=4

OpenStudy (anonymous):

please tell what is 7(3)x

OpenStudy (anonymous):

yes it is 7(3)^x....sorry i fell asleep =P

OpenStudy (anonymous):

are you sure about the question

OpenStudy (anonymous):

yeh i'm just clueless as to the answer!

OpenStudy (anonymous):

the first equation is 7(3)^(x-1)....the (-1) is included in the degree

OpenStudy (anonymous):

is the answer c if yes do what i said before h(x)g(x) put x=4

OpenStudy (anonymous):

are u not sure of the equation?

OpenStudy (anonymous):

what is your answer zupari

OpenStudy (anonymous):

Cant find an answer if the equation is wrong

OpenStudy (anonymous):

i think the method is correct

OpenStudy (anonymous):

I think I got it finally. h(x) = 4x x means current level right? So I f you are in the 4th level according to g(x) =7(3)^(x-1) g(x) = 189 h(x) = 4x h(x) = 4*4 =16 needed point = 189 * 16 =3024

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i dont know if the answer is correct or not

OpenStudy (anonymous):

Of course its correct

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

zupari you are from

OpenStudy (anonymous):

Sri lanka, where are you from?

OpenStudy (anonymous):

india

OpenStudy (anonymous):

Oh nearest neighbour

OpenStudy (anonymous):

ya i have to go know take care my friend

OpenStudy (anonymous):

Yeah you too